OpenStudy (anonymous):

The infinite sequence whose general term is an = 0.17n^2 - 1.02n + 6.67 where n=1,2,3,...,9 models the total operating costs, in millions of dollars, for a company from 1991 through 1999. find

OpenStudy (anonymous):

\[\sum_{i=1}^{5}ai\]

OpenStudy (anonymous):

i know this is probably simple but because it is in word problem form I can not figure out where to start

OpenStudy (klimenkov):

You have to find a sum \[\sum_{i=1}^5(0.17i^2-1.02i+6.67)\]

OpenStudy (anonymous):

so do I just place 1,2,3,4,5 in i and solve

OpenStudy (klimenkov):

Try to simplify it using \[\sum_{i=1}^ni^2=\frac{n(n+1)(2n+1)}6\]and \[\sum_{i=1}^ni=\frac{n(n+1)}2\]

OpenStudy (anonymous):

so n is 5 right

OpenStudy (anonymous):

or do I still do 1,2,3,4,5 as n and add the results altogether

OpenStudy (klimenkov):

You can do anything you like. Just sum all the addends in any way.

OpenStudy (anonymous):

I can't seem to get the right answer I got 35 but that is not one of my choices

OpenStudy (klimenkov):

Do you have an answer 27.4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have always hated word problems they seem to get the best of me

OpenStudy (anonymous):

thank you

OpenStudy (klimenkov):

No problems. Ask anything you like.